OpenStudy (anonymous):

What is the soulouition to the porportion? 9/x = 4/12

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

It may make the problem slightly easier to do to write it this way:\[9/x = 4/12\rightarrow \frac{ 9 }{ x }=\frac{ 12 }{ 4 }\] I'll bet you've seen equations like this before. How do you solve such equations?

OpenStudy (anonymous):

@hanner_B_nanner

OpenStudy (hanner_b_nanner):